A charge q is placed at the centre of the line joining two equal charges Q.Show that the system of three charges in equilibriumiq q=-Q/4 .

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
Draw a line OAB with O and B at the ends and A at the centre.
OA = AB
IF OB = x
Then OA = OB = x/2
        Force on Q placd at B due to point O(F1) = kQ2/x2
        Force on Q placd at B due to point A(F2)  = kQq/ x2/4
For the system be in equilibrium
F1+F2 = 0
kQ2/x2 +4 kQq/ x2= 0
Q+ 4Qq = 0
  • q = -Q/4
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)